# The Critical LYZ equation in Kähler Geometry

### v1

@@ 3.3. The complex Hessian estimate of Hou-Ma-Wu type
where we used $w_{i\bar{1},1}=w_{1\bar{1},i}$ since $d\chi=0$.
We prove
$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}\geq 0.$ (30)
This is true if $\lambda_{n}\geq 0$. If $\lambda_{n}<0$, by (16), we have
$\big|\frac{w_{n\bar{n},1}}{1+\lambda_{n}^{2}}\big|=\big|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\big|$
By the above equality and noting that $\lambda_{i}>0$ for $1\leq i\leq n-1$, we get
$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}=$ $\sum_{i=1}^{n-1}\frac{\lambda_{i}|w_{i\bar{i},1}|^{2}}{(1+\lambda_{i}^{2})^{2}}+\lambda_{n}\big|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\big|^{2}$
$\geq$ $\big(\sum_{i=1}^{n-1}\lambda_{i}^{-1}\big)^{-1}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}+\lambda_{n}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}$
$=$ $\frac{1+\lambda_{n}\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}{\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}$
$\geq$ $0,$
where in the last inequality we used
@@
Combining (3.3) with (30), we have
$F^{i\bar{i}}w_{i\bar{i},1\bar{1}}\geq\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}.$ (31)
Then we get
$\lambda_{1}^{-1}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}\geq$ $-\lambda_{1}^{-2}{(1+\lambda_{1}^{2})^{-1}}\big|w_{1\bar{1},1}|^{2}$
$-\sum_{i=2}^{n}\frac{|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}$
@@
We first estimate $\mathcal{I}$ and $\mathcal{II}$ by (25):
@@
$\geq$ $-\varphi^{\prime\prime}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}KF^{1\bar{1}}$ (33)
@@
$\geq$ $-2(1+\lambda_{1}^{2})^{-1}(\varphi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}-2(1+\lambda_{1}^{2})^{-1}(\psi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2}$
@@
where in the last inequality we used (27).
Next we estimate term $\mathcal{III}$ . Since $\lambda_{i}>0$ for $1\leq i\leq n-1$ by Lemma 2.2, we get
$\mathcal{III}\geq$ $\frac{\lambda_{n}|w_{1\bar{1},n}|^{2}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{n}^{2})},$
then by (25) and the Cauchy-Schwarz inequality, we obtain
@@
where we used (22). Let $\tau=\frac{1}{8A}$. We continue to estimate $\mathcal{III}$ by two cases.
Case 1: $|\lambda_{n}|\leq\tau\lambda_{1}$. In this case, we have
@@
Case 2: $|\lambda_{n}|\geq\tau\lambda_{1}$. By (27), we have $|u_{n\bar{n}}|^{2}=(\lambda_{n}-\chi_{n\bar{n}}-t)^{2}\geq\frac{\tau^{2}}{2}\lambda_{1}^{2}$ and we obtain
$4(\psi^{\prime})^{2}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}F^{n\bar{n}}|u_{n}|^{2}\leq$ $4AF^{n\bar{n}}K\leq\frac{1}{3}{\varphi^{\prime}}F^{n\bar{n}}u_{n\bar{n}}^{2}.$
Inserting the above inequality into (3.3), we have
@@
$\lambda_{1}^{-1}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}+\frac{3}{4}\varphi^{\prime\prime}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}+\psi^{\prime\prime}F^{i\bar{i}}|u_{i}|^{2}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}$
@@
where we used $u_{1\bar{1}}\geq\frac{\lambda_{1}}{2}$ which follows from (27).
Next we estimate $F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}$. By direct computation, we have
$F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}=F^{i\bar{i}}(u_{k\bar{i}}u_{\bar{k}i}+u_{ki}u_{\bar{k}\bar{i}})+F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}}).$ (40)
By covariant derivative formulae and (16), we have
$F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}})$
$\geq$ $-F^{i\bar{i}}(u_{i\bar{i}k}u_{\bar{k}}+u_{k}u_{i\bar{i}\bar{k}})-|Rm||\nabla u|^{2}\sum_{i}F^{i\bar{i}}$
$\geq$ $-2K(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}.$ (41)
@@
$F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}\geq F^{i\bar{i}}u_{i\bar{i}}^{2}-2K(|\nabla\chi|+|Rm|)\sum_{i=1}^{n}F^{i\bar{i}}.$ (42)
Inserting (28), (3.3) and (42) into (3.3), we obtain
$0\geq F^{i\bar{i}}Q_{i\bar{i}}\geq$ $\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}-3K\varphi^{\prime}(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}$
$\geq$ $\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}-2(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}},$ (43)
By Lemma 3.3 and noting that we have assumed $\underline{u}=0$, one of the following holds
$(i)\ -\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}\geq\delta_{0}\sum_{i=1}^{n}F^{i\bar{i}},\ \text{or}$
$(ii)\ F^{1\bar{1}}\geq\delta_{0}\sum_{i=1}^{n}F^{i\bar{i}}.$
We choose $C_{0}=6\delta_{0}^{-1}(\max|\nabla\chi|+|Rm|)$. If $(i)$ holds, by (3.3) and $\psi^{\prime}\geq C_{0}+1$, we get a contradiction. Then $(ii)$ holds and we obtain $|\lambda_{n}|\geq\frac{\delta_{0}}{2}\lambda_{1}$. Hence we have
$0\geq$ $\frac{\lambda_{1}^{2}}{4}F^{n\bar{n}}-3K(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}$
$\geq$ $\big(\frac{\lambda_{1}^{2}}{4}-3nK(|\nabla\chi|+|Rm|)\big)F^{n\bar{n}}.$
Therefore we get $\lambda_{1}\leq CK$ and $C$ is a uniform positive constant which is independent of $t$. ∎
4. Gradient estimates and the proof of the main theorem
@@ 4.1. Viscosity solutions
$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))+a_{0}\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\geq 0.$ (45)
@@
$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})+a_{0}\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})\geq 0.$
@@
Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible. As $v$ is a viscosity solution (and hence a supersolution), and $w+\epsilon|z|^{2}$ touches $v$ from above at $z_{0}$, the supersolution property yields:
$\sigma_{n-1}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})+a_{0}\sigma_{n}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})\leq 0.$ (46)
@@
$\sigma_{n-1}(w_{\alpha\bar{\beta}}(z_{0}))+a_{0}\sigma_{n}(w_{\alpha\bar{\beta}}(z_{0}))\leq-\epsilon^{n-1}<0.$ (47)
@@ 4.2. Weak solutions in the sense of currents
For the function $\tilde{w}(\tilde{z})=w(z)+|z_{n+1}|^{2}$ defined on $\mathbb{C}^{n+1}$, where $\tilde{z}=(z,z_{n+1})$, we define its mollification as
@@
    If $w$ is another subsolution, then $\frac{1}{2}(v+w)$ is a weak subsolution.
@@
$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\tilde{\varphi})(0,z_{0})\geq 0.$
@@ 4.3. A Liouville theorem to the equation $\sigma_{n}+\sigma_{n-1}=0$ in $\mathbb{C}^{n}$
$\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)+[v]_{\rho}(z)-\frac{7}{6}\bigr)$ $\geq v(x_{k}),$ (52)

### v2

@@ 3.3. The complex Hessian estimate of Hou-Ma-Wu type
where we have used $w_{i\bar{1},1}=w_{1\bar{1},i}$, which follows from $d\chi=0$.
We now prove the following inequality
$\sum_{i}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}\geq 0.$ (30)
This is clearly true if $\lambda_{n}\geq 0$. Assume now that $\lambda_{n}<0$. By (16), we have
$\Bigl|\frac{w_{n\bar{n},1}}{1+\lambda_{n}^{2}}\Bigr|=\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|.$
Using this relation and the fact that $\lambda_{i}>0$ for $1\leq i\leq n-1$, we obtain
$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}$ $=\sum_{i=1}^{n-1}\frac{\lambda_{i}|w_{i\bar{i},1}|^{2}}{(1+\lambda_{i}^{2})^{2}}+\lambda_{n}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
$\geq\Bigl(\sum_{i=1}^{n-1}\lambda_{i}^{-1}\Bigr)^{-1}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}+\lambda_{n}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
$=\frac{1+\lambda_{n}\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}{\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
$\geq 0,$
where the last inequality follows from
@@
Combining (3.3) with (30), we obtain
$\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}\geq\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}.$ (31)
Consequently,
$\lambda_{1}^{-1}\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}|w_{1\bar{1},i}|^{2}$
$\geq$ $-\lambda_{1}^{-2}(1+\lambda_{1}^{2})^{-1}|w_{1\bar{1},1}|^{2}-\sum_{i=2}^{n}\frac{|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}$
@@
We first estimate $\mathcal{I}$ and $\mathcal{II}$ using (25):
@@
$\geq-\varphi^{\prime\prime}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}KF^{1\bar{1}},$ (33)
@@
$\geq-2(1+\lambda_{1}^{2})^{-1}(\varphi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}$
$\quad-2(1+\lambda_{1}^{2})^{-1}(\psi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2}$
@@
where the last inequality follows from (27).
Next we estimate the term $\mathcal{III}$. Since $\lambda_{i}>0$ for $1\leq i\leq n-1$ by Lemma 2.2, we obtain
$\mathcal{III}\geq\frac{\lambda_{n}|w_{1\bar{1},n}|^{2}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{n}^{2})}.$
Then, using (25) and the Cauchy-Schwarz inequality, we have
@@
where we used (23).
Let $\tau=\frac{1}{8A}$. We proceed to estimate $\mathcal{III}$ in two cases.
Case 1: $|\lambda_{n}|\leq\tau\lambda_{1}$. In this case, we obtain
@@
Case 2: $|\lambda_{n}|\geq\tau\lambda_{1}$. By (27), we have $|u_{n\bar{n}}|^{2}=(\lambda_{n}-\chi_{n\bar{n}}-t)^{2}\geq\frac{\tau^{2}}{2}\lambda_{1}^{2}$, and consequently
$4(\psi^{\prime})^{2}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}F^{n\bar{n}}|u_{n}|^{2}\leq 4AF^{n\bar{n}}K\leq\frac{1}{3}\varphi^{\prime}F^{n\bar{n}}u_{n\bar{n}}^{2}.$
Inserting this inequality into (3.3) yields
@@
$\lambda_{1}^{-1}\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}|w_{1\bar{1},i}|^{2}$ (39)
$+\frac{3}{4}\varphi^{\prime\prime}\sum_{i}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}+\psi^{\prime\prime}\sum_{i}F^{i\bar{i}}|u_{i}|^{2}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}$
@@
where we used $u_{1\bar{1}}\geq\frac{\lambda_{1}}{2}$, which follows from (27).
Next we estimate $\sum\limits_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}$. By direct computation, we have
$\sum_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}=\sum_{i}F^{i\bar{i}}(u_{k\bar{i}}u_{\bar{k}i}+u_{ki}u_{\bar{k}\bar{i}})+\sum_{i,k}F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}}).$ (41)
Using the covariant derivative formulae and (16), we obtain
$\sum_{i,k}F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}})$
$\geq$ $-\sum_{i,k}F^{i\bar{i}}(u_{i\bar{i}k}u_{\bar{k}}+u_{k}u_{i\bar{i}\bar{k}})-|\text{Rm}||\nabla u|^{2}\sum_{i}F^{i\bar{i}}$
$\geq$ $-2K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (42)
@@
$\sum_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}\geq\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-2K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (43)
Inserting (28), (39) and (43) into (3.3), we obtain
$0\geq\sum_{i}F^{i\bar{i}}Q_{i\bar{i}}$ $\geq\psi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-3K\varphi^{\prime}(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}$
$\geq\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-2(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (44)
By Lemma 3.3 and noting that we have assumed $\underline{u}=0$, one of the following holds:
$(i)\ -\sum_{i}F^{i\bar{i}}u_{i\bar{i}}\geq\delta_{0}\sum_{i}F^{i\bar{i}},\quad\text{or}$
$(ii)\ F^{1\bar{1}}\geq\delta_{0}\sum_{i}F^{i\bar{i}}.$
We choose $C_{0}=6\delta_{0}^{-1}(\max|\nabla\chi|+|\text{Rm}|)$.
If (i) holds, then by (3.3) and $\psi^{\prime}\geq C_{0}+1$, we obtain a contradiction. Thus (ii) holds, and we get $|\lambda_{n}|\geq\frac{\delta_{0}}{2}\lambda_{1}$. Consequently, we have
$0$ $\geq\frac{\lambda_{1}^{2}}{4}F^{n\bar{n}}-3K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}$
$\geq\Bigl(\frac{\lambda_{1}^{2}}{4}-3nK(|\nabla\chi|+|\text{Rm}|)\Bigr)F^{n\bar{n}}.$
Therefore, $\lambda_{1}\leq CK$, where $C$ is a uniform positive constant independent of $t$. ∎
4. Gradient estimates and the proof of the main theorems
@@ 4.1. Viscosity solutions
$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))+\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\geq 0.$ (46)
@@
$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})+\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})\geq 0.$
@@
Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible. As $v$ is a viscosity solution (and hence a supersolution), by the definition of the supersolution, we have
$\sigma_{n-1}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})+\sigma_{n}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})\leq 0.$ (47)
@@
$\sigma_{n-1}(w_{\alpha\bar{\beta}}(z_{0}))+\sigma_{n}(w_{\alpha\bar{\beta}}(z_{0}))\leq-\epsilon^{n}<0.$ (48)
@@ 4.2. Weak solutions in the sense of currents
For the function $\tilde{w}(\tilde{z})=w(z)+|z_{n+1}|^{2}$ defined on $\mathbb{C}^{n+1}$, where $\tilde{z}=(z,z_{n+1})$, we define
@@
    If $w$ is a subsolution to the same equation, then $\frac{1}{2}(v+w)$ is also a subsolution.
@@
$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\tilde{\varphi})(z_{0},0)\geq 0.$
@@ 4.3. A Liouville theorem to the equation $\sigma_{n}+\sigma_{n-1}=0$ in $\mathbb{C}^{n}$
$\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(x_{k})+[v]_{\rho}(x_{k})-\frac{7}{6}\bigr)$ $\geq v(x_{k}),$ (53)